# What One Cannot, Two Can: Two-Layer Transformers Provably Represent Induction Heads on Any-Order Markov Chains

### v1

@@ D.2.1 Warmup: Second Order Markov Chain
[…]
@@ D.2.2 Any-Order Markov Chain
[…]
@@
Appendix E Stage 1 Lemmas
@@
Lemma 2 (Exension of Lemma G.1. in Nichani et al. [2024] to $k$th-order).
@@
The proof follows similarly to the proof of Lemma G.1 in Nichani et al. [2024]. We begin by bounding the difference inside the expectation.
@@
Lemma 3 (Extension of Lemma D.3 in Nichani et al. [2024]).
@@
The proof follows similarly to Nichani et al. [2024]. The gradient $G^{(1)}(A^{(1)},A^{(2)})_{i}$ can be expanded as
@@
Lemma 4 (Extension of Lemma D.5 in Nichani et al. [2024]: Bounding the time of convergence).
@@
The proof follows similarly to the proof of Lemma D.5 in Nichani et al. [2024]. Let $\mathcal{I}=\{1,\ldots,k\}$. Then, by Lemma 3, we have that throughout training,
@@ Appendix G Additional Lemmas
Lemma 7 (Lemma D.2 in Nichani et al. [2024]).
@@
Lemma 8 (Levin and Peres [2017]).
@@
Lemma 9 (Variance Lemma similar to Cor. F.6 in Nichani et al. [2024] ).
@@
Lemma 10 (Extension-1 to kth-order of Lemma G.2 in Nichani et al. [2024]).
@@
The proof is the same as the proof of Lemma G.2 in Nichani et al. [2024], where to control the variance term, we use 9. For completeness, we desribe the proof below:
@@
Lemma 11 (Extension-2 to kth-order of Lemma G.2 in Nichani et al. [2024]).
@@ H.1 Additional Experiments
In this section, we present additional attention map visualizations obtained by training a two-layer, single-head transformer on first-, second-, and third-order Markov chains, respectively. Attention maps from both the first and second layers are shown. For the first layer, we compute the average attention map over multiple sequences to understand the expected attention pattern. For the second layer, we display the attention map for a specific sequence. Additionally, we include a Pseudo Attention Map, which highlights the tokens that would ideally be attended to under a $k$-gram estimator. To assess alignment, we also plot the absolute difference between the actual attention map and the pseudo attention map. Note that for the first layer, we do not compute a pseudo attention map; hence, the pseudo attention and the corresponding absolute difference are not meaningful in that context.
@@ H.1.3 Third-Order Markov Chains
H.2 Experimental Details
@@
Architecture Based on the GPT-2 architecture as implemented in Pagliardini

### v2

@@ D.2.1 Warmup: Second Order Markov Chain
$\left.\frac{\mathrm{d}L(\theta)}{\mathrm{d}p_{1}}\right|_{i}=\left.\frac{\mathrm{d}L(\theta)}{\mathrm{d}p_{2}}\right|_{i}>\left.\frac{\mathrm{d}L(\theta)}{\mathrm{d}p_{k}}\right|_{i}\quad\forall\,k\in\{0,3,\ldots,n\}$
We denote $d_{g_{m,i}}=\left.\frac{\mathrm{d}L(\theta)}{\mathrm{d}p_{m}}\right|_{i}$. With this notation, we can express:
@@ D.2.2 Any-Order Markov Chain
$\left.\frac{\mathrm{d}L(\theta)}{\mathrm{d}p_{1}}\right|_{i}=\left.\frac{\mathrm{d}L(\theta)}{\mathrm{d}p_{2}}\right|_{i}\cdots=\left.\frac{\mathrm{d}L(\theta)}{\mathrm{d}p_{k}}\right|_{i}>\left.\frac{\mathrm{d}L(\theta)}{\mathrm{d}p_{q}}\right|_{i}\quad\forall\,q\in\{0,k+1,\ldots,n\}$
We denote $d_{g_{m,i}}=\left.\frac{\mathrm{d}L(\theta)}{\mathrm{d}p_{m}}\right|_{i}$. With this notation, we can express:
@@
Appendix E Training Dynamics: Stage One Lemmas
@@
Lemma 2 (Exension of Lemma G.1. in nichani2024how to $k$th-order).
@@
The proof follows similarly to the proof of Lemma G.1 in nichani2024how. We begin by bounding the difference inside the expectation.
@@
Lemma 3 (Extension of Lemma D.3 in nichani2024how).
@@
The proof follows similarly to nichani2024how. The gradient $G^{(1)}(A^{(1)},A^{(2)})_{i}$ can be expanded as
@@
Lemma 4 (Extension of Lemma D.5 in [nichani2024how]: Bounding the time of convergence).
@@
The proof follows similarly to the proof of Lemma D.5 in nichani2024how. Let $\mathcal{I}=\{1,\ldots,k\}$. Then, by Lemma 3, we have that throughout training,
@@ Appendix G Additional Lemmas
Lemma 7 (Lemma D.2 in nichani2024how).
@@
Lemma 8 (levin2017markov).
@@
Lemma 9 (Variance Lemma similar to Cor. F.6 in nichani2024how ).
@@
Lemma 10 (Extension-1 to kth-order of Lemma G.2 in nichani2024how).
@@
The proof is the same as the proof of Lemma G.2 in nichani2024how, where to control the variance term, we use 9. For completeness, we desribe the proof below:
@@
Lemma 11 (Extension-2 to kth-order of Lemma G.2 in nichani2024how).
@@ H.1 Additional Experiments
In this section, we present additional attention map visualizations obtained by training a two-layer, single-head transformer on first, second, and third-order Markov chains, respectively. Attention maps from both the first and second layers are shown. For the first layer, we compute the average attention map over multiple sequences to understand the expected attention pattern. For the second layer, we display the attention map for a specific sequence. Additionally, we include a Pseudo Attention Map, which highlights the tokens that would ideally be attended to under a $k$-gram estimator. To assess alignment, we also plot the absolute difference between the actual attention map and the pseudo attention map. Note that for the first layer, we do not compute a pseudo attention map; hence, the pseudo attention and the corresponding absolute difference are not meaningful in that context.
@@ H.1.3 Third-Order Markov Chains
H.2 Experiments on One Layer Transformers
To assess whether single-layer transformers can implement induction heads, we trained one-layer models on sequences generated from second-order Markov chains, using the same hyperparameter ranges as in our two-layer baselines.
Quantitative results.
As shown in Tab. 2, the gap between the true loss (computed from the ground-truth transition probabilities) and the loss achieved by the trained model remains significantly larger for one-layer transformers than for their two-layer counterparts. This difference persists across hyperparameter settings and indicates that deeper models capture the underlying structure more effectively.
Number of Layers Excess Cross-Entropy Loss
1 $0.131\pm 0.000$
2 $0.100\pm 0.003$
Table 2: Excess cross-entropy loss (loss obtained by the model subtracted with the true loss) for one and two-layer transformers trained on sequences from a second-order Markov chain.
Qualitative results.
Examination of the attention maps (Fig. 17, Fig. 18, Fig. 19) confirms this finding: one-layer transformers converge to a qualitatively different structure than the pseudo-attention map required for a $k$-gram estimator (see 1(b)). Specifically, attention mass concentrates along the two to three lower diagonals, rather than following the induction-like pattern.
These results are consistent with prior observations [olsson2022context] that one-layer transformers are insufficient for representing induction heads, while two-layer models succeed in doing so (albeit on different datasets).
Figure 17: Attention map corresponding to a randomly sampled sequence (denoted as sequence–1)
Figure 18: Attention map corresponding to a randomly sampled sequence (denoted as sequence–2)
Figure 19: Attention map corresponding to a randomly sampled sequence (denoted as sequence–3)
Figure 20: Attention map corresponding to a randomly sampled sequence (denoted as sequence–4)
H.3 Experiments on Noisy Sequences
We evaluate the robustness of our $k$-gram framework under non-ideal conditions by introducing stochastic noise into synthetic sequences generated from second-order Markov chains. We consider two noise processes: (i) random token substitution and (ii) perturbation of the transition dynamics. Models are two-layer, single-head transformers trained with three random seeds. Performance is reported as the difference between the model’s cross-entropy loss and the cross-entropy computed under the true transition probabilities over a large evaluation batch.
H.3.1 Random Token Substitution
$p_{n}$ Difference in Cross-Entropy Loss
0.0 $0.10\pm 0.003$
0.1 $0.15\pm 0.010$
0.25 $0.18\pm 0.002$
0.5 $0.21\pm 0.005$
Table 3: Cross-entropy loss difference under token substitution noise. Values are mean $\pm$ std. across three seeds.
We corrupt sequences at the token level with probability $p_{n}$. For each token, with probability $p_{n}$ we replace it with a uniformly sampled alternative; with probability $1-p_{n}$ it is left unchanged. The results can be found in Tab. 3. In the noiseless case ($p_{n}=0$), the learned second-layer attention map closely aligns with the pseudo-attention map (the optimal attention pattern for an ideal $k$-gram estimator). As $p_{n}$ increases, the alignment degrades and the loss gap widens.
H.3.2 Perturbation of Transition Dynamics
We next perturb the sequence dynamics directly. Let $P_{d}$ denote the true transition probability matrix. Before each prediction step we interpolate between $P_{d}$ and a randomly generated transition matrix $P_{r}$, defining
$P\;=\;(1-\alpha)\,P_{d}\;+\;\alpha\,P_{r},\qquad\alpha\in[0,1],$ (23)
where a new transition matrix $P_{r}$ is resampled at every time step. As in the previous setting, we measure the difference between the model’s cross-entropy and the true loss. Note that the results can be found in Tab. 4. The loss gap increases with $\alpha$. For $\alpha\leq 0.1$, attention maps remain close to the pseudo-attention pattern; at $\alpha=0.25$ and beyond, alignment deteriorates substantially. The model is more robust to this transition noise than to token-level corruption.
$\alpha$ Difference in Cross-Entropy Loss
0.0 $0.100\pm 0.003$
0.1 $0.101\pm 0.003$
0.25 $0.141\pm 0.005$
0.5 $0.181\pm 0.004$
Table 4: Cross-entropy loss difference under transition perturbations defined in Eq. (23). Values are mean $\pm$ std. across three seeds.
H.3.3 Discussion
These results complement our theory: when the data-generating process is close to Markovian, the learned attention structure approximates the pseudo-attention map; as noise increases, alignment degrades in tandem with cross-entropy performance. Overall, transformer-based estimators are sensitive to deviations from ideal $k$-gram dependencies, with greater robustness observed for transition perturbations than for token substitution.
H.4 Experiments on Bit Precision
Our bit-precision results are the same as those of [rajaraman2024transformers], which show that $\Omega(\log T+k)$ bits per parameter suffice for an additive error of $\mathcal{O}(1/T)$, where $T$ is the sequence length and $k$ the Markov order. To validate this, we quantize weights and activations of a model trained on a second-order Markov chain with sequence length 32. We then compute the Frobenius norm between the pseudo (optimal) attention map and the second-layer attention map of the quantized model across 10 random sequences (each from a different Markov transition kernel), and report the mean and variance:
Quantization (Bits) Frobenius Norm (Mean $\pm$ Std)
2 $4.6108\,\pm\,0.7190$
4 $4.5842\,\pm\,0.9826$
8 $3.8615\,\pm\,0.4590$
32 (No quantization) $3.7700\,\pm\,0.4725$
These results suggest that 8-bit quantization preserves the attention structure (as measured by the Frobenius norm) almost as well as full precision (32-bit), and aligns with the theoretical bounds.
H.5 Experimental Details
@@
Architecture Based on the GPT-2 architecture as implemented in [pagliardini-llm]
@@
Appendix I Parameter Breakdown and Comparison
In this section, we present a comprehensive comparison of parameter counts that explicitly accounts for all components of the models under consideration. Our analysis shows that our construction achieves greater parameter efficiency than the design of rajaraman2024transformers and also improves upon the approach proposed by nichani2024how. To make this concrete, we provide a full parameter breakdown for both our model and that of rajaraman2024transformers, detailing the contributions of embeddings, attention layers, feed-forward modules, and output projection.
Our Construction
The total parameter count of our construction is: $9(6S+3)^{2}+(6S+3)(2T+2S+9)$
Component Parameters
Token Embedding $(6S+3)\times S$
Attention Layer 1
Positional Encoding $(6S+3)\times T$
Query, Key, Value Projections $3(6S+3)^{2}$
MLP — (Three Layers)
Weights $3(6S+3)^{2}$
Biases $3(6S+3)$
Layer Norms $6(6S+3)$
Attention Layer 2
Positional Encoding $(6S+3)\times T$
Query, Key, Value Projections $3(6S+3)^{2}$
Output Projection $(6S+3)\times S$
Total $9(6S+3)^{2}+(6S+3)(2T+2S+9)$
Table 6: Parameter breakdown for our construction.
Construction in rajaraman2024transformers
The total parameter count of the construction in rajaraman2024transformers is: $15(6S+3)^{2}+(6S+3)(3T+2S+10)$
Component Parameters
Token Embedding $(6S+3)\times S$
Transformer Layer 1
Positional Encoding $(6S+3)\times T$
Query, Key, Value Projections $3(6S+3)^{2}$
MLP (2 layers) Weights $2(6S+3)^{2}$
MLP (2 layers) Biases $2(6S+3)$
Layer Norms $2(6S+3)$
Transformer Layer 2
Same as Transformer Layer 1: $(6S+3)\times T+5(6S+3)^{2}+4(6S+3)$
Transformer Layer 3
Positional Encoding $(6S+3)\times T$
Query, Key, Value Projections $3(6S+3)^{2}$
MLP (2 layers) Weights $2(6S+3)^{2}$
MLP (2 layers) Biases $2(6S+3)$
Layer Norms $2(6S+3)$
Output Projection $(6S+3)\times S$
Total $15(6S+3)^{2}+(6S+3)(3T+2S+10)$
Table 7: Parameter breakdown for the construction in rajaraman2024transformers.